OpenStudy (mayaal):

giving a medal!! find the exact value of the real number y. y=csc^-1(-1)

OpenStudy (anonymous):

same as solving \[\sin(x)=-1\] for \(x\)

OpenStudy (mayaal):

what does the -1 exponent on csc mean?

OpenStudy (anonymous):

inverse function

OpenStudy (mayaal):

but csc is not the inverse on sin,correct?

OpenStudy (anonymous):

\[\csc^{-1}(x)=y\iff csc(y)=x, -\frac{\pi}{2}\leq y\leq \frac{\pi}{2}\]

OpenStudy (anonymous):

on no inverse as in inverse function, not the reciprocal!!

OpenStudy (anonymous):

\[\csc(x)=\frac{1}{\sin(x)}\] but \[\sin^{-1}(x)\neq \csc(x)\] arcsine is the inverse function, not the reciprocal of the sine function

OpenStudy (mayaal):

so,what will we do after that?

OpenStudy (anonymous):

just like if \[f(x)=2x+1\] then \[f^{-1}(x)=\frac{x-1}{2}\] not \(\frac{1}{2x+1}\)

OpenStudy (mayaal):

ooooooh,u r a genius!

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

arccosecant is the inverse of cosecant you are being asked for the arccosecant of -1 that means you want a number whose cosecant is -1

OpenStudy (mayaal):

yes:)

OpenStudy (anonymous):

now you have no such button on your calculator so now we have to figure out what this really means

OpenStudy (anonymous):

since the cosecant is the reciprocal of the sine function (not the inverse, the reciprocal) if the cosecant of a number is -1, then the sine of it must also be -1

OpenStudy (mayaal):

yup

OpenStudy (mayaal):

so,do we look at the unit circle for the answer?

OpenStudy (anonymous):

yes

OpenStudy (mayaal):

is he answer 3pi/2?

OpenStudy (anonymous):

but on the unit circle there are an infinite number of numbers for which the sine is -1 one number is what you said, so it is tempting to say the answer is \(\frac{3\pi}{2}\) but that would be wrong

OpenStudy (mayaal):

do we want the general solutions,then?

OpenStudy (anonymous):

your answer has to lie in the interval from \(-\frac{\pi}{2}\) up to \(\frac{\pi}{2}\) and \(\frac{3\pi}{2}\) is not in that interval

OpenStudy (anonymous):

no, not a general solution arccosecant is supposed to be a well defined function, so we only want one number not a list of numbers

OpenStudy (mayaal):

do we find a coterminal angle?

OpenStudy (anonymous):

yes but don't think too hard

OpenStudy (mayaal):

ok,so: 3pi/2-2pi=-pi/2?

OpenStudy (anonymous):

yes, it should be \(-\frac{\pi}{2}\) look in your book to find the range of arccosecant it is a strange one

OpenStudy (mayaal):

the range of cosecant is:all real numbers?

OpenStudy (mayaal):

what is the range, @satellite73 ?

OpenStudy (anonymous):

hold on not the range of cosecant, the range of arccosecant

OpenStudy (mayaal):

oh,ok,i'm looking for it:)

OpenStudy (mayaal):

my book doesn't have it:(

OpenStudy (anonymous):

the range of cosecant is not all real numbers, it is \((-\infty, -1]\cup [1,\infty)\) the range of cosecant is \[(-\pi,-\frac{\pi}{2}]\cup (0,\frac{\pi}{2})\]

OpenStudy (anonymous):

sorry that second one should be the range of ARCCOSECANT

OpenStudy (mayaal):

thankyou,soooo much!

OpenStudy (anonymous):

yw